The Ajax Company can sell gidgets for 10.97 a piece. What does it cost the company to manifacture each gidget if it makes a profit of 8%.

i need a break down of why the answer is x=10.16;the manufactering cost is $10.16

Answer by checkley77(12844)

8%=8/100=.08
Cost of manufactuting the gidget=x.
x+.08x=selling price.
1.08x=selling price.
1.08x=10.97
x=10.97/1.08
x=10.16 cost of manufacturing the gidget.
--------------------------------
10.97/1.08=10.16 answer.